OpenStudy (anonymous):

how to integrate 1/(x^2+10*x+6)

OpenStudy (anonymous):

This is a repeat....partial fractions (or the tedious book method).

OpenStudy (anonymous):

\[\int{\frac{dx}{x^2 + 10x + 6}}\]

OpenStudy (anonymous):

i did this already. be happy to do it again, or let someone else do it

OpenStudy (anonymous):

partial fractions will get it. using arctanh will too but that is not intuitive

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quadratic_integral

OpenStudy (anonymous):

\[1/(x ^{2}+10x+25-19) =1/((x+5)^{2}-(\sqrt{19})^{2}\]\[(1/2*\sqrt{19})\log(\sqrt{19}+ x+5/\sqrt{19}-(x+5))\]hope u understood